# Scalable Parameter Design for Superconducting Quantum Circuits with Graph Neural Networks

### v2

@@ Abstract
Figure 2: Effectiveness, efficiency and scalability. These are the detailed versions of Fig. 1(i). a. The average crosstalk errors using different methods on SQECs with various structures and scales. b. The real-world running times of different algorithms. In both a and b, the horizontal axes represent the averaged qubit numbers of the test graphs, ranging from medium to large scales. Here, we test the SQECs with average qubit numbers from $31$ to $870$. The vertical axes correspond to the error and time cost respectively. All axes are expressed in logarithmic coordinates. The application (usage) of our GNNs-based designer is depicted in blue, the Snake algorithm is in orange, the direct application of gradient descent algorithm is in green, and the direct utilization of a gradient-free optimization algorithm is in purple. All the tests are performed using the same CPU and the same evaluator. Only the GNNs-based designer and the Snake algorithm are available for the large-scale graphs. In these figures, the GNNs-based designer demonstrates advantages in terms of effectiveness, efficiency and scalability.
@@
Discussion– This study initially demonstrates the advantages of applying GNNs to the hardware-level optimization of quantum chips, especially for large-scale ones. We apply two graph-based neural-network models named evaluator and designer to tackle the challenge of mitigating quantum crosstalk errors in large-scale SQECs. The evaluator can estimate crosstalk for large-scale SQECs containing thousands of qubits. And the designer outperforms traditional optimization algorithms in efficiency, effectiveness, and scalability. In the future research, additional error sources and variations between qubits may be taken into account, and the evaluator may be trained using experimental data, thus the practicality of the proposed algorithm can be further enhanced.
@@
Acknowledgments–We acknowledge J.-L. Long, Z. Wang, R.-B. Wu, G.-Z. Zhu, Y. Shang, T.-T. Chen, and G.-S. Liu for discussions. This work was support by Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300201).
Code Availability–The code will be open-soured after the official publication of the paper.
Contents in Supplementary Materials–In Supplementary Materials, we discuss the details of the algorithms and the SQECs based on several references [57, 58, 51, 52, 7, 8, 10, 36, 36, 59, 48, 53, 54, 37, 38, 50, 60, 61, 62, 63, 64, 65, 66, 67, 49, 28, 29, 30, 68], encompassing $8$ pages with $6$ images.

### v3

@@ Abstract
Figure 2: Effectiveness, efficiency and scalability. These are the detailed versions of Fig. 1(i). (a) The average crosstalk errors using different methods on SQECs with various structures and scales. (b) The real-world running times of different algorithms. All the tests are performed using the same CPU and the same evaluator. Only the GNNs-based designer and the Snake algorithm are available for the large-scale graphs.
Results of the designer–As shown in Fig. 1, during training the designer, medium-scale graphs $\mathcal{G}=(\mathcal{N},\mathcal{E})$ with nodes $\mathcal{N}$ and edges $\mathcal{E}$ are randomly generated and input into the designer to output a frequency assignment $\vec{\omega}$ for each graph. The trained evaluator estimates the crosstalk errors for a particular $\vec{\omega}$ on corresponding $\mathcal{G}$. Then the designer’s loss could be calculated,
$\mathcal{L}^{\rm designer}(\mathcal{G},\vec{\omega})=w_{X}\bar{E}_{R_{X}}+w_{Y}\bar{E}_{R_{Y}}+w_{XY}\bar{E}_{R_{XY}},$ (2)
where $\bar{E}_{O}$ is the average error for each target node under operation $O$ [53]. The default setting of the error weights $w_{O}$ is $w_{X}=w_{Y}=w_{XY}=1$. The randomly-generated training graphs contain around $54$, $103$ and $218$ nodes, and the training batch size is $288$. Approximately $700,000$ training steps were executed for about $2$ days with an RTX 4090 GPU. The trained designer can be readily applied to graphs of different structures and scales. We tested totally $30\times 6$ random graphs with average qubit numbers of $32$, $54$, $103$, $218$, $450$, and $870$. Other methods were also tested for comparison, including the direct application of a gradient-free optimizer, the direct utilization of gradient descent, and the Snake algorithm. We note that the trained evaluator is also necessary for the traditional methods, because conducting real experiments would require significant time and resources [28], and numerical simulation for large-scale SQECs is also challenging. All four methods employ the same trained evaluator on the same i9-14900KF CPU. Fig. 2(a) demonstrates that the well-trained GNNs-based designer achieves lower crosstalk errors than all other methods for all the graph scales. Additionally, the GNNs-based designer is exponentially faster than all other methods for all given tasks, as shown in Fig. 2(b). Especially for large-scale graphs where direct optimizations are not feasible, the GNNs-based designer achieves slightly more than half the errors of the Snake algorithm in several seconds, while the Snake algorithm would take dozens of minutes.
@@
Discussion– This study demonstrates the advantages of applying GNNs to the hardware-level optimization of quantum chips for the first time, especially for large-scale ones. We apply two graph-based neural-network models named evaluator and designer to tackle the challenge of mitigating quantum crosstalk errors in large-scale SQECs. The evaluator can estimate crosstalk for large-scale SQECs containing thousands of qubits. And the designer outperforms traditional optimization algorithms in efficiency, effectiveness, and scalability. In the future research, additional error sources and variations between qubits may be taken into account, and the evaluator may be trained using experimental data, thus the practicality of the proposed algorithm can be further enhanced.
@@
Acknowledgments–We acknowledge J.-L. Long, Z. Wang, R.-B. Wu, G.-Z. Zhu, Y. Shang, T.-T. Chen, and G.-S. Liu for discussions. This work was support by Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300200), and the National Natural Science Foundation of China with Grant No. 92365209.
Code Availability–The supporting code, data, and trained models for this study are openly available from GitHub [56].